4-(y-3) = x+3 1st equation

10+x = 2+(y-12)2nd equation
----
3


Answer by Alan3354(69443)

4-(y-3) = x+3 1st equation
10+x = 2+(y-12) 2nd equation
------------
1st eqn
4-y+3 = x+3
x+y=4
-------
2nd eqn
10+x = 2+y-12
x-y = -20
Add eqns 1 and 2
2x = -16
x = -8
So y = 12